OpenStudy (anonymous):

integral of 7cos^t(t) dt

OpenStudy (dumbcow):

\[\int\limits_{}^{}7 \cos t = 7 \sin t +C\] or is this \[\int\limits_{}^{}7 \cos^{t} t \] ?

OpenStudy (anonymous):

sorry wrote it wrong. integral of 7cos^7(t) dt

OpenStudy (dumbcow):

use reduction formula http://www.intmath.com/methods-integration/table-common-integrals.php look up cos^n and plug in 7 for n

OpenStudy (dumbcow):

Another option is to expand cos^7 \[7\int\limits_{?}^{?}\cos^{7}(t) = 7\int\limits_{?}^{?}\cos^{6}(t)\cos(t) =7\int\limits_{?}^{?}(\cos^{2}(t))^{3}\cos(t)=7\int\limits_{?}^{?}(1-\sin^{2}(t))^{3}\cos(t)\] The substitute u = sin t du = cos t \[\rightarrow 7\int\limits_{?}^{?}(1-u^{2})^{3} du\]